a particle of mass m and charge -qis moving around a charge +q2 along a circular path of radius r . find period of revolution of charge -q1?

Grade:12

1 Answers

Khimraj
5 years ago
centrifugal force = electrostatic force
mv2/r = k*q1*q2/r2
then v = \sqrt{k*q1*q2/m*r}
then T = 2\pir/v = 2\pi\sqrt{m*r/k*q1*q2}
Hope it clears.